# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ 5.3. Limitation and Future Works
  • develop metrics—both qualitative and on-chain quantitative—for assessing participation, emergent behaviour, and sociocultural impact;
  • integrate perspectives from art history, HCI, STS, economics, and legal studies to ground a theory of “protocolism” that distinguishes it from adjacent genres such as generative art and instruction-based conceptual art.
@@ 6. Conclusion
Our case study of Pak’s nine landmark protocol artworks reveals seven distinct characteristics of protocol art, helping advance both its definition and our understanding of this emerging genre. Protocol art remains an emerging creative domain that lacks a comprehensive theoretical framework. While ”protocolism” conceptualizes protocols as an artistic medium, the field needs more rigorous systematization of its concepts and aesthetics. Current discourse around blockchain-based art has emphasized market dynamics and hype over conceptual understanding. This rapid evolution has created a significant gap between practice and theory—one that requires careful critical reflection and theoretical development to properly ground the field.
@@ Appendix A Pak’s Protocol Artworks
This appendix contains collected information analyzing Pak’s protocol art through seven key lenses summarized and derived from ChatGPT-o3 Deep Research: (1) Concept, (2) System-mechanism dynamics, (3) Participatory interaction, (4) Artist’s control, (5) Collective emergent behavior, (6) Poetic meaning-making, and (7) Key statistics, followed by (8) Information.
@@ A.1. X (2020)
Figure 2. The Wave, Edition of 3, X by Pak (2020)
@@
Collectors engaged with X by acquiring the NFTs during a strictly limited timeframe, effectively collaborating in a one-day event rather than competing for scarce editions. Anyone could mint as many as desired within the 24-hour window, turning the drop into a collective performance synchronized in time. This format, then-unusual, invited a broad pool of participants to be part of the work’s creation – an approach that soon became a common practice in crypto art . The emphasis was on open participation: the art market’s typical frenzy for limited pieces was replaced by an inclusive but fleeting opportunity, underscoring the role of community presence and timing in the piece.
@@
After X was launched, Pak’s direct control was minimal – the rules were preset in the smart contract, and the outcome (how many tokens minted) was determined by the community within the allowed time. Pak’s role was chiefly in conceptualizing and coding the parameters; once the clock started, the process was autonomous and irreversible, with Pak acting more as instigator than controller. This ceding of control to the protocol is central to the work’s philosophy: the artist designed the playground but did not interfere with how many editions were ultimately created or who obtained them. The result is that X lives on as a completed event etched on-chain, with Pak’s influence embedded in the mechanism but not exerted in real-time during or after the sale.
Collective Emerging Behavior
X attracted a broad swath of NFT collectors due to its open nature, with every buyer during that day becoming part of the work’s narrative. The final mint counts effectively recorded the number of participants – a transparent metric of engagement – and in total, thousands of NFTs were minted, involving a large community. Rather than fostering a competitive race for a single token, X created a shared experience: collectors knew the availability was equal for all but also ticking away, which galvanized social media buzz and a sense of camaraderie among those “in the moment.” This collective temporal convergence – everyone minting in the same 24-hour span – became a form of emergent community behavior, reinforcing Pak’s notion that the artwork was as much about the crowd’s response as the digital objects themselves.
@@
X transforms the notion of an artwork into a time-bound performance, offering a poetic commentary on how value and meaning in digital art are shaped by temporal and social parameters. The piece demonstrates that an NFT’s significance can lie not just in its image but in the circumstances of its creation – here, a collective moment that cannot be replicated. By foregrounding the protocol (the 24-hour clock) over the visual content, Pak underscored themes of ephemerality and community: the fact that X could only come into being through synchronized action imbues it with a narrative of shared presence. Ultimately, X invites reflection on an alternate mode of art valuation: one where scarcity is a function of time and collective presence rather than material limitation, thereby poetically aligning the artwork with the ethos of a decentralized, time-sensitive digital culture.
@@
Released 27 August 2020 on Nifty Gateway , the X collection ultimately comprised 15 artworks with a total of 63 editions minted (across 13 open works and 2 one-of-ones) . Over 24 hours, collectors minted as few as 3 and at most 8 copies per open-edition piece, totaling only 61–64 pieces—far fewer than the theoretical infinity, underscoring the experiment’s success. The highest known sale from X was the unique The Void, which was auctioned to a single collector (price undisclosed publicly), while aggregate primary sales for X exceeded six figures in USD terms (each open edition priced in the low hundreds of dollars). This series marked Pak’s first major NFT drop, establishing the artist’s reputation for novel mechanics in crypto art .
@@ A.2. The Title (2020)
Figure 3. The Title by Pak (2021)
@@
The Title (December 2020) is a conceptual NFT series that interrogates the nature of value and ownership in crypto art. Pak created nine virtually identical digital objects – all a rotating translucent cube – but each NFT was given a different title and method of distribution (e, 2021). By selling the same image at different prices and under labels like The Cheap, The Expensive, The Unsold, etc., Pak demonstrated that the perceived value of an NFT lies not in the visual content but in contextual factors such as scarcity, nomenclature, and narrative. The artwork’s core concept is essentially a self-referential critique of the NFT medium: The Title asks what it is we truly buy when we purchase an NFT, positing that the token’s title and associated story can override the image in defining its essence.
@@
Under the hood, The Title utilized a clever technical setup: all nine NFTs pointed to the exact same IPFS-hosted media file of the spinning cube. Despite this shared imagery, the smart contract distinguished each token by metadata – notably the title and edition info – which conferred each piece’s unique identity and implied value. This meant that ownership of any one token was effectively ownership of the same underlying image, forcing a paradox: collectors held different tokens but “shared” the art content. The dynamics of this system laid bare the role of metadata and token provenance in NFT valuation; questions arose such as whether the holder of “The Expensive” had more claim to the art than the holder of “The Cheap,” even though the visual asset was identical. By structuring the NFTs in this way, Pak’s smart contract became a philosophical device, highlighting how blockchain metadata and distribution rules can generate distinct meanings and hierarchies around an otherwise fungible digital image.
@@
Collectors engaged with The Title through a gamified multi-format sale. Each of the nine NFTs was obtained via a different mechanism – some were open editions (available to anyone for a limited time), others were limited editions at set prices, one was sold via blind auction, one via a standard auction, and one (“The Gift”) was not sold at all but given away (e, 2021). This design meant that the audience had to navigate various modes of acquisition, effectively turning the act of collecting into part of the artwork’s performative narrative. By strategically involving buyers in auctions, giveaways, and timed drops, Pak cultivated a community-wide discourse and excitement around the project. The process of obtaining the pieces became a participatory spectacle of its own, reinforcing the idea that the artwork extends beyond the image to include the protocols and social interactions of its distribution.
@@
Pak’s control over The Title after release was primarily conceptual and pre-programmed. Once the pieces were distributed through their various channels, Pak did not intervene in the tokens’ existence or content – in fact, by anchoring all tokens to a single immutable IPFS image, he relinquished any ability to differentiate or alter the visual component for one token without affecting them all. The only lever of control exerted was at inception: defining the titles and scarcity of each edition, thereby scripting the value narrative in advance. After launch, the market of collectors took over; the artist’s role shifted to observer of the unfolding debate and trade. This limited post-launch control aligns with the work’s intention: it left collectors to ascribe meaning and value among themselves, illustrating Pak’s point that the community’s perceptions and the immutable smart contract rules (not the artist’s hand) ultimately govern an NFT’s fate.
Collective Emerging Behavior
@@
The poetic impact of The Title lies in its elegant exposure of art-world conventions transposed to the blockchain. By echoing Yves Klein’s 1957 exhibition of identical blue paintings sold at different prices(Blu, 2020), Pak placed NFT culture in dialogue with historical avant-garde inquiries into what art really is. The Title makes viewers ponder why one iteration of a digital cube should carry more prestige or value than another. It reveals that the poetry of the piece comes from the questions it raises: What is ownership? What do we value in art – the image, the story, or the token that confers bragging rights? The work’s broader significance is in demonstrating that blockchain art can critique itself; it’s simultaneously a satire and a celebration of the cryptographic medium, showing how scarcity, community perception, and narrative alchemy collectively create meaning in digital art. In sum, The Title stands as a meta-artwork that is reflexive about its own value structure, engaging the art discourse on authenticity, authorship, and the economics of the intangible.
@@
Launched 5 January 2021 on Nifty Gateway, The Title comprised 9 works: The Cheap (192 editions at \$499 each), The Expensive (8 editions at \$10,000 each), The Unsold (1 edition, priced so high it remained unsold), The Blind (3 editions via silent auction), The Flipper (99 editions first-come at a low price), The Lucky (3 editions first-come, price \$1 each – effectively a lottery of speed), Copy and Paste (each 1/1 auctions, paired conceptually), and The Gift (3 editions given free to early collectors). Total sales exceeded \$200,000 across the series. Critically, all nine tokens reference one image file , illustrating Pak’s point. Sotheby’s later noted this collection as a breakthrough in conceptual NFT art, aligning Pak with avant-garde strategies of value interrogation.
@@ A.3. The Fungible
Figure 4. The Fungible by Pak (2021)
@@
The Fungible (April 2021) was an ambitious collection that probed the definition of value in the context of digital art and fungibility. Created in partnership with Sotheby’s, this project’s very title is a play on the idea of fungible vs. non-fungible value: Pak offered an array of NFTs that collectively challenged the boundaries between unique art objects and interchangeable units. The concept centered on asking “what does value mean, and from where does it derive authority?” (Fun, 2021) — a question that the collection explored by introducing novel mechanics (like token merging and dynamic editions) which blurred the line between a singular artwork and a set of exchangeable pieces. In essence, The Fungible served as a critical examination of how scarcity and abundance intersect on the blockchain, turning the sale itself into an inquiry about why we value art: is it the content, the context of rarity, or the creative token logic behind it?
@@
The Fungible deployed innovative smart-contract mechanics that made collecting an interactive process. Core to the system were the Open Edition “Cubes,” which were not traditional 1-of-n editions but a kind of modular token: buyers could purchase any number of identical cubes, and the contract would then deliver NFTs reflecting the quantities accumulated. For instance, if a collector purchased five cubes, they would receive an NFT depicting a cluster of five cubes; if they purchased fifty, they’d receive a single NFT showing fifty cubes — a graduated series that cleverly visualized ownership as a spectrum. This algorithmic distribution turned fungible purchases into bespoke non-fungible artworks. Additionally, Pak introduced dynamic one-of-one NFTs like The Switch, coded such that its owner could decide to “flip” the artwork’s state once at a moment of their choosing (an irreversible transformation built into the contract). Another piece, The Pixel, consisted of a single gray pixel, emphasizing minimalism and the idea that even the smallest digital unit could hold enormous value in the right context. These mechanics – custom distribution, owner-triggered change, and extreme minimalism – showcased the versatility of smart contracts, engendering a set of artworks whose form and existence were defined by procedural rules and collector interaction, not just by static visuals.
@@
The Fungible’s drop was designed as a multi-day, multi-modal event that actively involved the audience in creating and discovering value. Over a 48-hour period, collectors could participate in Open Edition sales where they bought “fungible cubes” in whatever quantity they wished during the sale period (Fun, 2021), effectively letting each participant construct their own edition. Meanwhile, high-stakes auctions for one-of-a-kind pieces (The Switch and The Pixel) ran in parallel, drawing competitive bidding. Pak further engaged the community with puzzles and special rewards: for example, he gifted 30 unique NFTs to notable community “builders” and awarded four more NFTs (called “Equilibrium”) to participants who met specific criteria like solving a puzzle or tweeting Pak’s hashtag to the largest audience (202, 2021b). This layered approach meant that thousands of people joined at various levels – from casual buyers to dedicated players – making the sale itself feel like a collective happening. Every participant, whether buying a single cube or vying in an auction, became part of the narrative fabric that The Fungible wove around the notion of value.
@@
Once The Fungible launched, Pak largely let the programmed systems and collectors drive the outcomes. The rules for minting cubes and the parameters for pieces like The Switch were all predetermined, meaning that after initiating the sale, Pak did not manually influence how many cubes were sold or how the artworks evolved. In fact, The Switch exemplified a deliberate transfer of artistic agency: the moment of its transformation was controlled by the collector, not the artist (Fun, 2021). Similarly, the open edition ran on its own schedule – Pak set the prices and time windows for each cube sale, but could not alter the open-edition once it was live. By structuring the project this way, Pak’s role became one of orchestrator rather than micromanager. He created the conditions and constraints, then ceded control to the market and the code, a choice consistent with the work’s exploration of decentralized value creation. After the sale, the artworks existed independently on the blockchain (e.g., The Pixel perpetually remaining just one pixel owned by its auction winner), requiring no further intervention from the artist.
@@
The Fungible attracted an unprecedented scale of participation for an NFT art drop at the time. Over 3,000 individuals took part in the open edition cube sale alone, resulting in a total of 23,598 cubes sold across three short selling periods and 6,156 unique NFTs minted from the open edition process. The event’s gamified structure (with leaderboards for top cube buyers, surprise rewards, and the allure of auctions) encouraged a spirited community response. Buyers strategized over how many cubes to purchase to attain higher-tier cube NFTs, while others collaborated or competed for the special rewards. Sotheby’s reported record-breaking engagement on Nifty Gateway during the sale, including the most bids ever placed on an NFT auction item up to that point. This collective excitement transcended typical art-world transactions; it felt more akin to a massive multiplayer event in which the audience’s size and enthusiasm were integral to the work’s impact. In the aftermath, collectors and observers alike debated whether the frenzy and competitive “game” of The Fungible were integral to its value, or if those incentives would fade, directly engaging with the very question the artwork posed about how value is sustained in digital art.
@@
Beyond its clever mechanics and record-breaking sales, The Fungible resonated as a poetic statement on the fluid nature of value in the digital age. By literalizing the idea of fungibility – letting collectors accumulate and assemble value units (cubes) – Pak drew attention to the transactional element of art without forfeiting aesthetic intrigue. The collection as a whole becomes an allegory: the cubes symbolize the building blocks of value, and their conversion into unique artworks symbolizes the alchemy by which markets and perception turn the mundane into the precious. The open question The Fungible leaves us with is whether art’s value is intrinsic or constructed: when a single pixel sells for \$1.36 million and thousands of identical cubes become rare sculptures through a smart contract, one is compelled to acknowledge the role of consensus, context, and collective belief in creating worth. In the broader digital art discourse, The Fungible is seen as a work that married participatory art with commentary on economics, ultimately reminding us in a poignant way that even on the blockchain, value is a story we all collectively tell.

### v2

@@ 5.3. Limitations and Future Work
  • develop metrics—both qualitative and on-chain quantitative—for assessing participation, emergent behavior, and sociocultural impact;
  • integrate perspectives from art history, HCI, STS, economics, ethics (Calvo, 2024), and legal studies to ground a theory of “protocolism” that distinguishes it from adjacent genres such as generative art and instruction-based conceptual art.
@@ 6. Conclusion
Our case study of Pak’s nine landmark protocol artworks reveals seven distinct characteristics of protocol art, helping advance both its definition and our understanding of this emerging genre. Protocol art remains an emerging creative domain that lacks a comprehensive theoretical framework. While “protocolism” conceptualizes protocols as an artistic medium, the field needs more rigorous systematization of its concepts and aesthetics. Current discourse around blockchain-based art has emphasized market dynamics and hype over conceptual understanding (Poposki, 2024a). This rapid evolution has created a significant gap between practice and theory—one that requires careful critical reflection and theoretical development to properly ground the field.
@@ Appendix A Pak’s Protocol Artworks
This appendix contains collected information analyzing Pak’s protocol art through eight analytical lenses summarized and derived from ChatGPT-o3 Deep Research: (1) Concept, (2) System-mechanism dynamics, (3) Participatory interaction, (4) Artist’s control, (5) Collective emergent behavior, (6) Poetic meaning-making, (7) Key statistics, and (8) Annotation.
@@ A.1. X (2020)
Refer to caption A 3D-rendered sculptural form composed of layered horizontal slices arranged into an ellipsoidal shell, rendered in grayscale with soft lighting against a black background, bisected by a thin horizontal line.
Figure 2. The Wave, edition of 3, from Pak’s X collection (2020). Thirteen of the 15 pieces in X shared a time-based open-edition format where scarcity was determined by a 24-hour minting window rather than a fixed supply; the remaining two were one-of-one auctions.
@@
Collectors engaged with X by acquiring the NFTs during a strictly limited timeframe, effectively collaborating in a one-day event rather than competing for scarce editions. Anyone could mint as many as desired within the 24-hour window, turning the drop into a collective performance synchronized in time. This format, then-unusual, invited a broad pool of participants to be part of the work’s creation—an approach that soon became a common practice in crypto art. The emphasis was on open participation: the art market’s typical frenzy for limited pieces was replaced by an inclusive but fleeting opportunity, underscoring the role of community presence and timing in the piece.
@@
After X was launched, Pak’s direct control was minimal—the rules were preset in the smart contract, and the outcome (how many tokens minted) was determined by the community within the allowed time. Pak’s role was chiefly in conceptualizing and coding the parameters; once the clock started, the process was autonomous and irreversible, with Pak acting more as instigator than controller. This ceding of control to the protocol is central to the work’s philosophy: the artist designed the playground but did not interfere with how many editions were ultimately created or who obtained them. The result is that X lives on as a completed event etched on-chain, with Pak’s influence embedded in the mechanism but not exerted in real-time during or after the sale.
Collective Emergent Behavior
X attracted a broad swath of NFT collectors due to its open nature, with every buyer during that day becoming part of the work’s narrative. The final mint counts effectively recorded the number of participants—a transparent metric of engagement—and in total, thousands of NFTs were minted, involving a large community. Rather than fostering a competitive race for a single token, X created a shared experience: collectors knew the availability was equal for all but also ticking away, which galvanized social media buzz and a sense of camaraderie among those “in the moment.” This collective temporal convergence—everyone minting in the same 24-hour span—became a form of emergent community behavior, reinforcing Pak’s notion that the artwork was as much about the crowd’s response as the digital objects themselves.
@@
X transforms the notion of an artwork into a time-bound performance, offering a poetic commentary on how value and meaning in digital art are shaped by temporal and social parameters. The piece demonstrates that an NFT’s significance can lie not just in its image but in the circumstances of its creation—here, a collective moment that cannot be replicated. By foregrounding the protocol (the 24-hour clock) over the visual content, Pak underscored themes of ephemerality and community: the fact that X could only come into being through synchronized action imbues it with a narrative of shared presence. Ultimately, X invites reflection on an alternate mode of art valuation: one where scarcity is a function of time and collective presence rather than material limitation, thereby poetically aligning the artwork with the ethos of a decentralized, time-sensitive digital culture.
@@
Released 28 August 2020 on Nifty Gateway, the X collection ultimately comprised 15 artworks with a total of 63 editions minted (across 13 open works and 2 one-of-ones). Over 24 hours, collectors minted as few as 3 and at most 8 copies per open-edition piece, totaling only 63 pieces—far fewer than the theoretical infinity, underscoring the experiment’s success. The highest known sale from X was the unique The Void, which was auctioned to a single collector (price undisclosed publicly), while aggregate primary sales for X exceeded six figures in USD terms (each open edition priced in the low hundreds of dollars). This series marked Pak’s first major NFT drop, establishing the artist’s reputation for novel mechanics in crypto art.
Annotation
X is the foundational case for C1 (system-centric composition) and C4 (temporal dynamism): the artwork has no fixed edition size—its supply is a function of collective behavior within a 24-hour window, making the protocol’s time constraint, rather than any visual property, the locus of artistic meaning. What makes this particularly significant for digital art is the inversion of the scarcity paradigm: while artists can trivially copy digital files, market demand cannot be copied, and Pak exploits this asymmetry by substituting a temporal mechanism for the conventional predetermination of edition count. Scarcity in X is not declared by the artist but discovered through market-driven behavior—a limited window generates a limited edition without any authorial fiat over supply. This absence of predetermination is conceptually radical: it demonstrates that the scarcity of digital artwork need not be imposed from above but can emerge from the structure of participation itself. X thus initiates C5 (economy-driven engagement) by turning the act of purchasing into the generative act itself, carrying a poetic meaning of market dynamism in which economically driven engagement literally constitutes the work. Notably, X shows minimal C2 (autonomous governance) and no C7 (composability): it is a self-contained event with no token mechanics that bridge to later works. As Pak’s earliest NFT project, X establishes the core intuition—scarcity as a protocol parameter rather than an artist’s decree—that all subsequent works elaborate.
@@ A.2. The Title (2021)
Refer to caption A screenshot of the Nifty Gateway marketplace showing a three-by-three grid of nine NFT listings, each displaying the same rotating translucent 3D cube but with different titles and prices: The Cheap (499 dollars), The Expensive (10,000 dollars), The Unsold (1,000,000 dollars, zero editions minted), The Lucky (1 dollar), The Flipper (1 dollar), The Blind (silent auction), The Gift (not for sale), Copy (17,888 dollars, sold out), and Paste (17,777 dollars, sold out).
Figure 3. Marketplace listing for The Title by Pak (2021). All nine NFTs share an identical rotating cube visual; only the on-chain title, price, and distribution method distinguish each piece—foregrounding naming and context over visual content as determinants of value.
@@
The Title (January 2021) is a conceptual NFT series that interrogates the nature of value and ownership in crypto art. Pak created nine virtually identical digital objects—all a rotating translucent cube—but each NFT was given a different title and method of distribution (e, 2021). By selling the same image at different prices and under labels like The Cheap, The Expensive, The Unsold, etc., Pak demonstrated that the perceived value of an NFT lies not in the visual content but in contextual factors such as scarcity, nomenclature, and narrative. The artwork’s core concept is essentially a self-referential critique of the NFT medium: The Title asks what it is we truly buy when we purchase an NFT, positing that the token’s title and associated story can override the image in defining its essence.
@@
Under the hood, The Title utilized a clever technical setup: all nine NFTs pointed to the exact same IPFS-hosted media file of the spinning cube. Despite this shared imagery, the smart contract distinguished each token by metadata—notably the title and edition info—which conferred each piece’s unique identity and implied value. This meant that ownership of any one token was effectively ownership of the same underlying image, forcing a paradox: collectors held different tokens but “shared” the art content. The dynamics of this system laid bare the role of metadata and token provenance in NFT valuation; questions arose such as whether the holder of “The Expensive” had more claim to the art than the holder of “The Cheap,” even though the visual asset was identical. By structuring the NFTs in this way, Pak’s smart contract became a philosophical device, highlighting how blockchain metadata and distribution rules can generate distinct meanings and hierarchies around an otherwise fungible digital image.
@@
Collectors engaged with The Title through a gamified multi-format sale. Each of the nine NFTs was obtained via a different mechanism—some were open editions (available to anyone for a limited time), others were limited editions at set prices, one was sold via blind auction, one via a standard auction, and one (“The Gift”) was not sold at all but given away (e, 2021). This design meant that the audience had to navigate various modes of acquisition, effectively turning the act of collecting into part of the artwork’s performative narrative. By strategically involving buyers in auctions, giveaways, and timed drops, Pak cultivated a community-wide discourse and excitement around the project. The process of obtaining the pieces became a participatory spectacle of its own, reinforcing the idea that the artwork extends beyond the image to include the protocols and social interactions of its distribution.
@@
Pak’s control over The Title after release was primarily conceptual and pre-programmed. Once the pieces were distributed through their various channels, Pak did not intervene in the tokens’ existence or content—in fact, by anchoring all tokens to a single immutable IPFS image, Pak relinquished any ability to differentiate or alter the visual component for one token without affecting them all. The only lever of control exerted was at inception: defining the titles and scarcity of each edition, thereby scripting the value narrative in advance. After launch, the market of collectors took over; the artist’s role shifted to observer of the unfolding debate and trade. This limited post-launch control aligns with the work’s intention: it left collectors to ascribe meaning and value among themselves, illustrating Pak’s point that the community’s perceptions and the immutable smart contract rules (not the artist’s hand) ultimately govern an NFT’s fate.
Collective Emergent Behavior
@@
The poetic impact of The Title lies in its elegant exposure of art-world conventions transposed to the blockchain. By echoing Yves Klein’s 1957 exhibition of identical blue paintings sold at different prices (ArtWizard, 2020), Pak placed NFT culture in dialogue with historical avant-garde inquiries into what art really is. The Title makes viewers ponder why one iteration of a digital cube should carry more prestige or value than another. It reveals that the poetry of the piece comes from the questions it raises: What is ownership? What do we value in art—the image, the story, or the token that confers bragging rights? The work’s broader significance is in demonstrating that blockchain art can critique itself; it’s simultaneously a satire and a celebration of the cryptographic medium, showing how scarcity, community perception, and narrative alchemy collectively create meaning in digital art. In sum, The Title stands as a meta-artwork that is reflexive about its own value structure, engaging the art discourse on authenticity, authorship, and the economics of the intangible.
@@
Launched 6 January 2021 on Nifty Gateway, The Title comprised 9 works: The Cheap (192 editions at \$499 each), The Expensive (8 editions at \$10,000 each), The Unsold (1 edition, priced so high it remained unsold), The Blind (3 editions via silent auction), The Flipper (99 editions first-come at a low price), The Lucky (3 editions first-come, price \$1 each—effectively a lottery of speed), Copy and Paste (each 1/1 auctions, paired conceptually), and The Gift (3 editions given free to early collectors). Total sales exceeded \$200,000 across the series. Critically, all nine tokens reference one image file, illustrating Pak’s point. Sotheby’s later noted this collection as a breakthrough in conceptual NFT art, aligning Pak with avant-garde strategies of value interrogation.
Annotation
The Title is the purest expression of C5 (economy-driven engagement) and C6 (poetic message embedding) in the portfolio: by selling identical images at different prices under different names, Pak isolates price and metadata as the sole carriers of meaning, forcing participants to confront the question of whether value resides in the visual object or in the narrative apparatus that frames it. The satirical force of the piece lies in its provocation: since a smart contract requires a unique identifier for each entry, the work asks whether it is the name itself that holds primary value—whether the title, not the content, is what collectors truly acquire. This interrogation strikes at the heart of non-fungibility: even identical images, once assigned different names within the contract, become formally distinct, non-fungible things, revealing that the ontological boundary between fungible and non-fungible may hinge on something as thin as a string of metadata. The work thus exemplifies C1 (system-centric composition) at its most distilled, since the “artwork” is the distribution structure and naming convention itself rather than the rotating cube that all nine tokens share. The Title is, in this sense, a work of “protocol art” in the strictest definition: the contract is the piece, and its poetic message emerges from the nature of the contract’s requirements. However, The Title is an outlier for C3 (distributed agency): collectors cannot alter the work or contribute content—they can only choose which token to acquire. This constraint is itself instructive: it reveals that communal authorship is not essential for protocol art to function as critique, but its absence limits the emergent social dynamics observed in later projects.
@@ A.3. The Fungible
Refer to caption A screenshot of the Nifty Gateway marketplace showing a two-by-four grid of open-edition NFT listings. Each listing displays a different quantity of translucent 3D cubes arranged in grid formations against black backgrounds: A Cube (1 cube), Five Cubes, Ten Cubes, Twenty Cubes, Fifty Cubes, Hundred Cubes, Five Hundred Cubes, and Thousand Cubes. All are labeled Open Edition.
Figure 4. Open-edition tiers of The Fungible by Pak (2021). Collectors purchased cubes in varying quantities; the smart contract then minted unique NFTs whose visual composition reflected the number of cubes accumulated, turning fungible purchases into bespoke non-fungible artworks.
@@
The Fungible (April 2021) was an ambitious collection that probed the definition of value in the context of digital art and fungibility. Created in partnership with Sotheby’s, this project’s very title is a play on the idea of fungible vs. non-fungible value: Pak offered an array of NFTs that collectively challenged the boundaries between unique art objects and interchangeable units. The concept centered on asking “what does value mean, and from where does it derive authority?” (Sotheby’s, 2021)—a question that the collection explored by introducing novel mechanics (like token merging and dynamic editions) which blurred the line between a singular artwork and a set of exchangeable pieces. In essence, The Fungible served as a critical examination of how scarcity and abundance intersect on the blockchain, turning the sale itself into an inquiry about why we value art: is it the content, the context of rarity, or the creative token logic behind it?
@@
The Fungible deployed innovative smart-contract mechanics that made collecting an interactive process. Core to the system were the Open Edition “Cubes,” which were not traditional 1-of-n editions but a kind of modular token: buyers could purchase any number of identical cubes, and the contract would then deliver NFTs reflecting the quantities accumulated. For instance, if a collector purchased five cubes, they would receive an NFT depicting a cluster of five cubes; if they purchased fifty, they’d receive a single NFT showing fifty cubes—a graduated series that cleverly visualized ownership as a spectrum. This algorithmic distribution turned fungible purchases into bespoke non-fungible artworks. Additionally, Pak introduced dynamic one-of-one NFTs like The Switch, coded such that its owner could decide to “flip” the artwork’s state once at a moment of their choosing (an irreversible transformation built into the contract). Another piece, The Pixel, consisted of a single gray pixel, emphasizing minimalism and the idea that even the smallest digital unit could hold enormous value in the right context. These mechanics—custom distribution, owner-triggered change, and extreme minimalism—showcased the versatility of smart contracts, engendering a set of artworks whose form and existence were defined by procedural rules and collector interaction, not just by static visuals.
@@
The Fungible’s drop was designed as a multi-day, multi-modal event that actively involved the audience in creating and discovering value. Over a three-day period, collectors could participate in Open Edition sales where they bought “fungible cubes” in whatever quantity they wished during the sale period (Sotheby’s, 2021), effectively letting each participant construct their own edition. Meanwhile, high-stakes auctions for one-of-a-kind pieces (The Switch and The Pixel) ran in parallel, drawing competitive bidding. Pak further engaged the community with puzzles and special rewards: for example, he gifted 30 unique NFTs to notable community “builders” and awarded four more NFTs (called “Equilibrium”) to participants who met specific criteria like solving a puzzle or tweeting Pak’s hashtag to the largest audience (.ART, 2021). This layered approach meant that thousands of people joined at various levels—from casual buyers to dedicated players—making the sale itself feel like a collective happening. Every participant, whether buying a single cube or vying in an auction, became part of the narrative fabric that The Fungible wove around the notion of value.
@@
Once The Fungible launched, Pak largely let the programmed systems and collectors drive the outcomes. The rules for minting cubes and the parameters for pieces like The Switch were all predetermined, meaning that after initiating the sale, Pak did not manually influence how many cubes were sold or how the artworks evolved. In fact, The Switch exemplified a deliberate transfer of artistic agency: the moment of its transformation was controlled by the collector, not the artist (Sotheby’s, 2021). Similarly, the open edition ran on its own schedule—Pak set the prices and time windows for each cube sale, but could not alter the open-edition once it was live. By structuring the project this way, Pak’s role became one of orchestrator rather than micromanager. He created the conditions and constraints, then ceded control to the market and the code, a choice consistent with the work’s exploration of decentralized value creation. After the sale, the artworks existed independently on the blockchain (e.g., The Pixel perpetually remaining just one pixel owned by its auction winner), requiring no further intervention from the artist.
@@
The Fungible attracted an unprecedented scale of participation for an NFT art drop at the time. Over 3,000 individuals took part in the open edition cube sale alone, resulting in a total of 23,598 cubes sold across three short selling periods and 6,165 unique NFTs minted from the open edition process. The event’s gamified structure (with leaderboards for top cube buyers, surprise rewards, and the allure of auctions) encouraged a spirited community response. Buyers strategized over how many cubes to purchase to attain higher-tier cube NFTs, while others collaborated or competed for the special rewards. Sotheby’s reported record-breaking engagement on Nifty Gateway during the sale, including the most bids ever placed on an NFT auction item up to that point. This collective excitement transcended typical art-world transactions; it felt more akin to a massive multiplayer event in which the audience’s size and enthusiasm were integral to the work’s impact. In the aftermath, collectors and observers alike debated whether the frenzy and competitive “game” of The Fungible were integral to its value, or if those incentives would fade, directly engaging with the very question the artwork posed about how value is sustained in digital art.
@@
Beyond its clever mechanics and record-breaking sales, The Fungible resonated as a poetic statement on the fluid nature of value in the digital age. By literalizing the idea of fungibility—letting collectors accumulate and assemble value units (cubes)—Pak drew attention to the transactional element of art without forfeiting aesthetic intrigue. The collection as a whole becomes an allegory: the cubes symbolize the building blocks of value, and their conversion into unique artworks symbolizes the alchemy by which markets and perception turn the mundane into the precious. The open question The Fungible leaves us with is whether art’s value is intrinsic or constructed: when a single pixel sells for \$1.36 million and thousands of identical cubes become rare sculptures through a smart contract, one is compelled to acknowledge the role of consensus, context, and collective belief in creating worth. In the broader digital art discourse, The Fungible is seen as a work that married participatory art with commentary on economics, ultimately reminding us in a poignant way that even on the blockchain, value is a story we all collectively tell.